OpenStudy (maheshmeghwal9):

Calculate the electric field strength near a long charged wire, carrying charge per unit length "q".

OpenStudy (maheshmeghwal9):

@experimentX Plz hlp:)

OpenStudy (experimentx):

make use of gauss law.

OpenStudy (experimentx):

this is one of the earliest standard problem you study on electrostat ... probably you should google.

OpenStudy (maheshmeghwal9):

actually its answer is also some thing different it is \[E = \frac{1}{4 \pi \epsilon_0}\frac{2q}{r}\]

OpenStudy (maheshmeghwal9):

from where did this "2" come?

OpenStudy (experimentx):

I told you google it ... try to understand it ... and if you don't understand ask me ... deriving this is quite long.

OpenStudy (maheshmeghwal9):

ok:)

OpenStudy (maheshmeghwal9):

yeah from gauss law we gt this but how to get \[E = \frac{1}{4 \pi \epsilon_0}\frac{2q}{r}\]this?

OpenStudy (maheshmeghwal9):

this is the answer in my book i found nothing @ google about this:\

OpenStudy (maheshmeghwal9):

i mean from gauss law we get \[\frac{\lambda}{2 \pi r \epsilon_0}\]but how to get this \[E=\frac{1}{4 \pi \epsilon_0}\frac{2q}{r}\]

OpenStudy (maheshmeghwal9):

?????

OpenStudy (maheshmeghwal9):

oh ok i gt it thanx:D

OpenStudy (experimentx):

hold on

OpenStudy (maheshmeghwal9):

ok!

OpenStudy (maheshmeghwal9):

i seriously gt it

OpenStudy (maheshmeghwal9):

\[E=\frac{1}{4 \pi \epsilon_0}\frac{2q}{r}=\frac{q}{2 \pi r \epsilon_0}= \frac{\lambda}{2 \pi r \epsilon _0}.\] Since charge per unit length {lambda} is q.

OpenStudy (experimentx):

oh ... nice

OpenStudy (maheshmeghwal9):

:)